How to calculate period in magnetic field?


1
Expert's answer
2020-07-29T10:30:52-0400

If we have a charged particle in a magnetic field, then the period of its rotation can be determined using Newton's second law


"T=2\pi R/v"


"F=ma\to qvB\sin\alpha=m\frac{v^2\cdot \sin^2\alpha}{R}" "\to"


"v=\frac{qBR}{m\sin\alpha}"


Finally


"T=\frac{2\pi Rm\sin\alpha}{qBR}=\frac{2\pi m\sin\alpha}{qB}." Answer